# Exploring Dynamic Transformer for Efficient Object Tracking

### v1

@@ Abstract
The speed-precision trade-off is a critical problem for visual object tracking which usually requires low latency and deployment on constrained resources. Existing solutions for efficient tracking mainly focus on adopting light-weight backbones or modules, which nevertheless come at the cost of a sacrifice in precision. In this paper, inspired by dynamic network routing, we propose DyTrack, a dynamic transformer framework for efficient tracking. Real-world tracking scenarios exhibit diverse levels of complexity. We argue that a simple network is sufficient for easy frames in video sequences, while more computation could be assigned to difficult ones. DyTrack automatically learns to configure proper reasoning routes for various inputs, gaining better utilization of the available computational budget. Thus, it can achieve higher performance with the same running speed. We formulate instance-specific tracking as a sequential decision problem and attach terminating branches to intermediate layers of the entire model. Especially, to fully utilize the computations, we introduce the feature recycling mechanism to reuse the outputs of predecessors. Furthermore, a target-aware self-distillation strategy is designed to enhance the discriminating capabilities of early predictions by effectively mimicking the representation pattern of the deep model. Extensive experiments on multiple benchmarks demonstrate that DyTrack achieves promising speed-precision trade-offs with only a single model. For instance, DyTrack obtains 64.9% AUC on LaSOT with a speed of 256 $fps$.
@@ I Introduction
Visual object tracking aims to continuously locate an arbitrary target object in a video sequence. The astounding performance of current algorithms [4, 5, 3, 6] is obtained by deep and complex network architectures [7, 8, 9]. In particular, the successful application of the transformer [10] has pushed tracking performance to a new state-of-the-art. Despite top-notch performance, these trackers come at the expense of increased latency cost and, therefore, have limitations such as deployment on resource-constrained devices. They fall into “Cannikin’s Law” where higher burden structures are designed to handle a small number of complex cases. While in real-world object tracking, efficiency and performance should be both important.
Some studies [2, 11, 12, 13] dedicated to improving tracking efficiency have emerged. ECO [2] proposes a factorized convolution operator with lower linear complexity to reduce the number of parameters in DCF [14] model. LightTrack [12] utilizes the power of neural architecture search (NAS) to find lightweight network modules (i.e., backbone and head). These modules form a lightweight tracker, with fewer floating point operations per second (FLOPs) and parameters. Recently, Borsuk et al.  [13] present a fast siamese tracker consisting of a compact feature extraction block and a fusion module. These methods achieve satisfactory efficiency by dedicated designing or searching lightweight architectures or modules. However, suffering from the “speed or precision dilemma”, these efficient trackers lean towards speed, thus partially lacking the capabilities to handle complex and corner cases.
@@ II Related Work
Visual Object Tracking. Siamese network [21] and discriminative correlation filter (DCF) [14] dominate the field of visual object tracking over the past decades. Tracking performance is constantly being refreshed with ongoing efforts. SiamRPN [5] introduces the RPN [22] from detection into Siamese tracking framework, it gains accurate bounding boxes by offline training with large-scale image pairs from ILSVRC [23] and Youtube-BB [24]. Then, SiamRNP++[25] adopts deeper ResNet [8] for feature extraction, further enhancing the accuracy. Online trackers [26, 27, 2, 11, 28] obtain a clearer discriminative capability to interpret target and background by target-aware model fine-tuning. Recently, many trackers [3, 6, 29, 30, 31] introduce transformer as their backbones [9, 32] and transformer-based template-search integration modules [10]. TransT [3] proposes ego-context and cross-feature modules based on self-attention and cross-attention, respectively. More recently, many works [6, 29, 33, 30, 31] take transformer backbone for joint feature extraction and interaction. MixFormer [29] designs a backbone based on CVT [34] and OSTrack [6] uses an MAE [35] pre-trained ViT to joint extract the cues from the template and search region. GRM [31] designs a token division module to select search tokens to interact with the template tokens for alleviating target-background confusion, and more complex modeling relationships are learned. However, most works focus on the pursuit of higher performance while ignoring efficiency which should also be critical in practical applications.
Efficient Tracking. In real-world applications, tracking algorithms usually require high efficiency for diverse deployment purposes, such as unmanned vehicles and UAV vision. Some studies have explored approaches to achieve efficient tracking. ECO [2] introduces a factorized convolution operator to reduce the computational complexity in DCF models. It has a high inference speed, but the performance is inferior to meet the current tracking requirement. To meet the requirements of resource-limited platforms for model FLOPs and parameters, LightTrack [12] uses NAS [36] to search for a tracker with a lightweight backbone and prediction head. While in practice, FLOPs and the parameter number do not directly measure the actual speed of the model  [37]. More recently, FEAR [13] employs depth-wise separable convolutions and designs compact feature extraction and fusion blocks, and E.T.Track [38] proposes a single instance-level attention layer to build a transformer-based tracker. They both attain high energy efficiency. However, like high-performance trackers but at another extreme, these methods favor efficiency in the efficiency-precision trade-off. They remain constrained in dealing with some complex and corner cases. Based on SiamFC [21], EAST [39] achieves adaptive tracking with cascaded CNN features. Despite achieving a promising speedup, it needs to learn policies by complex reinforcement learning and has to train different models for different precision-speed trade-offs which lacks flexibility and utility. In contrast, we try to address the above concerns from the perspective that we formulate the speed-precision problem in a dynamic transformer network by instance-specific routing.
@@ III-A Overview
where $Comp$ represents the feature reuse manner (detailed descriptions in Section III-B). We keep the shape and dimension of $\bm{h}_{k}^{c}$ the same as $\bm{h}_{k}$. Note that $\pi_{1}$ only receives $\bm{h}_{1}$ as the input. Then the features are fed into an adapter block to learn a paradigm shift for discriminating the target object. The adapter is composed of several transformer layers. After a projection operation, the generated search tokens are candidates for inputs of the box prediction head. The IoU token, is fed into a 3-layer MLP to estimate the current frame’s IoU score which is the exiting criterion.
Figure 5: Diagrams of four feature reuse schemas.
@@ III-B Feature Recycling
Figure 6: Illustration of the target-aware self-distillation. Early features are learned with additional supervision from teacher from the deep layer. Imitation attention re-weight the student features to valuable regions in terms of teacher features.
@@ III-C Target-aware Self-distillation
DyTrack has multiple exiting routes to execute box prediction. For deeper layers, it is usually easier to learn high-level semantic representation [47]. Hence, the earlier the output node is, the harder an accurate tracking branch can be learned. To bridge this gap, we propose to encourage the shallow and middle branches to mimic the response pattern of the deep layer. Specifically, we introduce a target-aware self-distillation training strategy whose process is shown in Figure 6. In DyTrack, the deepest feature representation $\bm{f}_{K}$ acts as the teacher of all previous features $\{\bm{f}_{k}\}_{k=1}^{K-1}$, and multiple prediction branches distributed in various depths of the network enable conditions for self-distillation learning. It may not be optimal to directly distill all the features. Intuitively, the target response region is more important. To further promote the distillation, we introduce an imitation attention module $\mathcal{G}_{k}$ to re-weight the learned features of different regions, where $k\in[1,K-1]$. Note that, at the inference, these imitation attention modules will be removed without bringing additional computations. Specifically, $\mathcal{G}_{k}$ receives two inputs which are student tokens $\bm{f}_{k}$ from hidden branches and teacher tokens $\bm{f}_{K}$ from finally reasoning route. After respective reshape and transform (convolutional and linear layers) operations, $\bm{f}_{K}$ is projected to two groups of spatial attention maps $\bm{Att}_{s}$ and channel attention vectors $\bm{Att}_{c}$. These learned attention acts on the student features to obtain the target-aware feature representation $\bm{f}_{k}^{{}^{\prime}}$. Finally, we utilize the cosine similarity as the imitation measurement between the teacher representation $\bm{f}_{K}$ and the re-weighted student representation $\bm{f}_{k}^{{}^{\prime}}$ as:
@@
Without extra teachers, the target-aware self-distillation process are trained end-to-end.
@@ IV-A Implementation Details
Training Details. We employ the training splits of GOT-10k [54], TrackingNet [55], COCO2017 [66], and LaSOT [1] as our training data. For the evaluation of GOT-10k, we only use the training split of GOT-10k to follow the official requirements. We sample from the training sequence to generate the image pairs. For COCO2017 which is an imageg dataset, we execute different data augmentations on the same image the obtain the image pairs. The search region and template size are $4^{2}$ and $2^{2}$ the area of the object, which are resized to 256$\times$256 and 128$\times$128 patches before fed into the model. We train the model in two stages: we first train the backbone and the final prediction branch, then we include the other exiting branches for joint training. The training configurations of the two stages are the same. AdamW [67] is employed as our optimizer, with a weight decay of 1e-4. The learning rate of the backbone is set to 1e-5, and the remaining modules are set to 1e-4. We use 2 Nvidia A100 GPUs to train our models. In each stage, the model is trained for 300 epochs with a batch size of 64. Each epoch contains 60,000 sampling pairs. The learning rate decreases by a factor of 10 after 240 epochs.
@@ IV-B State-of-the-art Comparisons
IV-C Exploration Studies
Design choices for feature reuse. We introduce a feature recycling mechanism that reduces the waste and redundancy of computations by feature reuse. Hereby, we compare four variants of reuse manner as described in Section III-B. The corresponding exploratory experiments are shown in Table III. As shown in experiment #2, without feature reuse, the performance declines significantly, especially at the early and intermediate exits. It indicates that computations from early decisioner networks can be exploited to facilitate better box prediction in our dynamic transformer structure. Experiments #3, #1, #4, and #5 in Table III represent employing the feature reuse schemas (a), (b), (c), and (d) in Figure 5. We can find that residual integration can not bring improvements. This may be because the decisioner network recieves the input features from the encoder, so it is more suitable for learning direct representations of the target rather than its residuals. The concatenation operation also shows faintly helpful. We find the gated summation schema bring improvements compared with #2. But they introduce additional computations which reduce the efficiency. Therefore, we adopt the schema (b) which is a summation operation before being sent into decisioner networks, and we find this simple operation performs the best.
@@ VI Conclusion
This work proposes DyTrack, a dynamic transformer tracking framework for efficient object tracking. Different with existing efficient tracking methods that focus on designing lightweight module, DyTrack works with an instance-specific reasoning fashion that learns to assign various inference routes for different frames, reducing the computations for easy scenarios. We further introduce feature recycling and target-aware self-distillation to facilitate efficiency and precision. Extensive experiments demonstrate that DyTrack achieves a superior speed-precision trade-off compared with both high-performance and high-speed methods.

### v2

@@ Abstract
The speed-precision trade-off is a critical problem in visual object tracking, as it typically requires low latency and is deployed on resource-constrained platforms. Existing solutions for efficient tracking primarily focus on lightweight backbones or modules, which, however, come at a sacrifice in precision. In this paper, inspired by dynamic network routing, we propose DyTrack, a dynamic transformer framework for efficient tracking. Real-world tracking scenarios exhibit varying levels of complexity. We argue that a simple network is sufficient for easy video frames, while more computational resources should be assigned to difficult ones. DyTrack automatically learns to configure proper reasoning routes for different inputs, thereby improving the utilization of the available computational budget and achieving higher performance at the same running speed. We formulate instance-specific tracking as a sequential decision problem and incorporate terminating branches to intermediate layers of the model. Furthermore, we propose a feature recycling mechanism to maximize computational efficiency by reusing the outputs of predecessors. Additionally, a target-aware self-distillation strategy is designed to enhance the discriminating capabilities of early-stage predictions by mimicking the representation patterns of the deep model. Extensive experiments demonstrate that DyTrack achieves promising speed-precision trade-offs with only a single model. For instance, DyTrack obtains 64.9% AUC on LaSOT with a speed of 256 $fps$.
@@ I Introduction
Visual object tracking aims to continuously locate an arbitrary target object in a video sequence. The impressive performance of recent tracking algorithms [1, 2, 3, 4, 5, 6] is largely attributed to sophisticated network architectures [7, 8, 9]. In particular, the successful integration of the transformer [10] has significantly advanced tracking performance, setting new state-of-the-arts. Despite top-notch performance, these trackers come at the expense of increased latency cost and, therefore, have limitations such as deployment on resource-constrained devices. They fall into “Cannikin’s Law” where higher burden structures are designed to handle a small number of complex cases. While in real-world object tracking, efficiency and performance should be both important.
Some studies [11, 12, 13, 14, 15] dedicated to improving tracking efficiency have emerged. ECO [11] proposes a factorized convolution operator with lower linear complexity to reduce the number of parameters in DCF [16] model. LightTrack [13] utilizes the power of neural architecture search (NAS) to find lightweight network modules (i.e., backbone and head). These modules form a lightweight tracker, with fewer floating point operations per second (FLOPs) and parameters. Recently, Borsuk et al.  [14] present a fast siamese tracker consisting of a compact feature extraction block and a fusion module. These methods achieve satisfactory efficiency by dedicated designing or searching lightweight architectures or modules. However, suffering from the “speed or precision dilemma”, these efficient trackers lean towards speed, thus partially lacking the capabilities to handle complex and corner cases.
@@ II Related Work
Visual Object Tracking. Siamese network [24] and discriminative correlation filter (DCF) [16] dominate the field of visual object tracking over the past decades. Tracking performance is constantly being refreshed with ongoing efforts. SiamRPN [1] introduces the RPN [25] from detection into Siamese tracking framework, it gains accurate bounding boxes by offline training with large-scale image pairs from ILSVRC [26] and Youtube-BB [27]. Then, SiamRNP++[28] adopts deeper ResNet [8] for feature extraction, further enhancing the accuracy. Online trackers [11, 29, 12, 30, 31] obtain a clearer discriminative capability to interpret target and background by target-aware model fine-tuning. For example, Zhang et al.  [31] propose a attention-driven memory network for online tracking, in which an memory updater is designed to promote the model reliability. Recently, many trackers [3, 5, 6, 32, 33] introduce transformer as their backbones [9, 34] and transformer-based template-search integration modules [10]. TransT [3] proposes ego-context and cross-feature modules based on self-attention and cross-attention, respectively. More recently, many works [5, 6, 35, 36, 32, 33] take transformer backbone for joint feature extraction and interaction. MixFormer [6] designs a backbone based on CVT [37] and OSTrack [5] uses an MAE [38] pre-trained ViT to joint extract the cues from the template and search region. GRM [33] designs a token division module to select search tokens to interact with the template tokens for alleviating target-background confusion, and more complex modeling relationships are learned. From the perspective of training data, Li et al.  [39] propose to synthesize sufficient samples by a crop-transform-paste operation, boosting the performance of various supervised trackers. However, most works focus on the pursuit of higher performance while ignoring efficiency which should also be critical in practical applications.
Efficient Tracking. In real-world applications, tracking algorithms usually require high efficiency for diverse deployment purposes, such as unmanned vehicles and UAV vision. Some studies have explored approaches to achieve efficient tracking. ECO [11] introduces a factorized convolution operator to reduce the computational complexity in DCF models. It has a high inference speed, but the performance is inferior to meet the current tracking requirement. To meet the requirements of resource-limited platforms for model FLOPs and parameters, LightTrack [13] uses NAS [40] to search for a tracker with a lightweight backbone and prediction head. While in practice, FLOPs and the parameter number do not directly measure the actual speed of the model  [41]. Ussa et al.  [42] propose a real-time hybrid neuromorphic tracker that aggregates clues from event cameras for maintaining robust tracking in occluded scenarios. More recently, FEAR [14] employs depth-wise separable convolutions and designs compact feature extraction and fusion blocks, and E.T.Track [43] proposes a single instance-level attention layer to build a transformer-based tracker. They both attain high energy efficiency. However, like high-performance trackers but at another extreme, these methods favor efficiency in the efficiency-precision trade-off. They remain constrained in dealing with some complex and corner cases. LiteTrack [44] employs asynchronous feature extraction and interaction, and reduces redundant computation through the proposed encoder layer pruning. MixFormerV2 [45] proposes a fully transformer-based tracking framework that eliminates dense convolutional operations and complex score heads, achieving real-time speed on CPU platform. To use large-stride downsampling hierarchical backbone for tracking, HiT [46] incorporates the high-level information of deep features into the shallow large-resolution features. ABTrack [47] designs adaptively bypassing transformer and introduce a pruning method to speed up tracker inference. For efficient multi-modal tracking, Zhang et al. [48] introduce a cross-modality distillation framework that transfers effective knowledge from a powerful two-stream tracker to a compact single-stream tracker. These approaches improve efficiency by designing lighter architectures or utilizing pruning and distillation techniques. However, they are fixed-speed trackers tailored to specific platforms. When adapting to different speed-precision trade-offs, they require re-training or even re-designing the model. Based on SiamFC [24], EAST [49] achieves adaptive tracking with cascaded CNN features. Despite achieving a promising speedup, it needs to learn policies by complex reinforcement learning and has to train different models for different precision-speed trade-offs which lacks flexibility and utility. In contrast, we try to address the above concerns from the perspective that we formulate the speed-precision problem in a dynamic transformer network by instance-specific routing.
@@ III-A Overview
where $Comp$ represents the feature reuse manner (detailed descriptions in Section III-B). We keep the shape and dimension of $\bm{h}_{k}^{c}$ the same as $\bm{h}_{k}$. Note that $\pi_{1}$ only receives $\bm{h}_{1}$ as the input. Then the features are fed into an adapter block to learn a paradigm shift for discriminating the target object. The adapter is composed of several transformer layers. After a linear projection operation, the generated search tokens serve as candidates for the inputs to the box prediction head. Following previous studies [61, 62], the IoU token, is fed into a 3-layer MLP to estimate the current frame’s IoU score which is the exiting criterion.
@@ III-B Feature Recycling
Figure 5: Diagrams of four feature reuse schemas. The output of the score head is omitted for better readability.
@@ III-C Target-aware Self-distillation
DyTrack has multiple exiting routes to execute box prediction. For deeper layers, it is usually easier to learn high-level semantic representation [57]. Hence, the earlier the output node is, the harder an accurate tracking branch can be learned. To bridge this gap, we propose to encourage the shallow and middle branches to mimic the response pattern of the deep layer. Specifically, we introduce a target-aware self-distillation training strategy whose process is shown in Figure 6. In DyTrack, the deepest feature representation $\bm{f}_{K}$ acts as the teacher of all previous features $\{\bm{f}_{k}\}_{k=1}^{K-1}$, and multiple prediction branches distributed in various depths of the network enable conditions for self-distillation learning. It may not be optimal to equally distill all the features. Intuitively, the object response regions are more critical than the background regions. To further promote the distillation, we introduce an imitation attention module $\mathcal{G}_{k}$ to re-weight the learned features of different regions, where $k\in[1,K-1]$. The imitation attention mechanism is achieved by re-weighting the student features, which is implemented through spatial and channel attention derived from the teacher features. In this way, the student features can focus more effectively on the key regions and channels during the distillation process. Specifically, $\mathcal{G}_{k}$ receives two inputs which are student tokens $\bm{f}_{k}$ from hidden branches and teacher tokens $\bm{f}_{K}$ from finally reasoning route. After respective reshape and transform (convolutional and linear layers) operations, $\bm{f}_{K}$ is projected to two groups of spatial attention maps $\bm{Att}_{s}$ and channel attention vectors $\bm{Att}_{c}$. Specifically, the transform operation including a 1$\times$1 convolution layer and a linear projection layer. To obtain $\bm{Att}_{c}$, we additionally employ a $softmax$ operation in the spatial dimension of the features. These learned attention acts on the student features to obtain the target-aware feature representation $\bm{f}_{k}^{{}^{\prime}}$. Finally, we utilize the cosine similarity as the imitation measurement between the teacher representation $\bm{f}_{K}$ and the re-weighted student representation $\bm{f}_{k}^{{}^{\prime}}$ as:
@@
Without extra teachers, the target-aware self-distillation process are trained end-to-end. And note that, at the inference, the imitation attention modules will be removed without bringing additional computations.
Figure 6: Illustration of the target-aware self-distillation. Early features are learned with additional supervision from teacher from the deep layer. Imitation attention re-weight the student features to valuable regions in terms of teacher features. We omit the output of the score head for better readability.
@@ IV-A Implementation Details
Training Details. We employ the training splits of GOT-10k [66], TrackingNet [67], COCO [78], and LaSOT [17] as our training data. For the evaluation of GOT-10k, we only use the training split of GOT-10k to follow the official requirements. We sample from the training sequence to generate the image pairs. For COCO which is an imageg dataset, we execute different data augmentations on the same image to obtain the image pairs. The search region and template size are $4^{2}$ and $2^{2}$ the area of the object, which are resized to 256$\times$256 and 128$\times$128 patches before fed into the model. We train the model in two stages: we first train the backbone and the final prediction branch, then we include the other exiting branches for joint training. The training configurations of the two stages are the same. AdamW [79] is employed as our optimizer, with a weight decay of 1e-4. The learning rate of the backbone is set to 1e-5, and the remaining modules are set to 1e-4. We use 2 Nvidia A100 GPUs to train our models. In each stage, the model is trained for 300 epochs with a batch size of 64. Each epoch contains 60,000 sampling pairs. The learning rate decreases by a factor of 10 after 240 epochs.
@@ IV-B State-of-the-art Comparisons
IV-C Exploration Studies
Design choices for feature reuse. We introduce a feature recycling mechanism that reduces the waste and redundancy of computations by feature reuse. Hereby, we compare four variants of reuse manner as described in Section III-B. The corresponding exploratory experiments are shown in Table III. As shown in experiment #2, without feature reuse, the performance declines significantly, especially at the early and intermediate exits. It indicates that computations from early decisioner networks can be exploited to facilitate better box prediction in our dynamic transformer structure. Experiments #3, #1, #4, and #5 in Table III represent employing the feature reuse schemas (a), (b), (c), and (d) in Figure 5. We can find that residual integration can not bring improvements. This may be because the decisioner network recieves the input features from the encoder, so it is more suitable for learning direct representations of the target rather than its residuals. And applying element-wise summation to the input features allows the recycled features to undergo deeper representations, thereby providing more discriminative cues. The concatenation operation also proves to be slightly beneficial. Its performance is inferior to that of (b), primarily because (b) leverages the spatial alignment between recycled features and current features, enhancing critical spatial information while avoiding the additional computational cost introduced by the increased number of tokens after concatenation. We observe that the gated summation scheme brings improvements compared to #2. However, it does not achieve promising performance gains compared to schema (b), likely because the recycled features exhibit no significant redundancy. Moreover, the additional computations introduced by the gate operation reduce the efficiency. Therefore, we adopt schema (b), which applies a summation operation before sending features into the Decisioner networks. This simple operation is found to deliver the best performance.
@@ IV-C Exploration Studies
Figure 11: Evaluation of models of different sizes on test subsets with varying difficulty levels.
Performance analysis on varying difficulty levels. To further analyze the effect of model size on different difficulty levels, we divide the test sequences (from the LaSOT benchmark) into 5 difficulty levels using a small model. The small model is the E-1 model of DyTrack, processes inputs through the simplest path and directly outputs predictions from the 1st exit. By sorting the predicted results’ IoU scores across all sequences, we uniformly divide the sequences into 5 subsets (each containing 56 sequences), representing 5 difficulty levels. The IoU scores of the E-1 model on the test subsets are shown in Fig. 11 (a). For sequences with the lowest difficulty level (L-0), the smallest E-1 model achieves an average IoU of 87.7%. However, as the difficulty level increases, the IoU scores start to decline, reaching only 29.7% on the most challenging sequences at the L-4 difficulty level. When using larger (i.e., more complex) models, the performance consistently improves across all difficulty levels. A noticeable trend is that larger models yield significantly greater improvements on subsets with higher difficulty levels. For instance, the E-2 model achieves only a 1.2% IoU improvement over the base E-1 model on L-0 sequences but demonstrates a substantial 5.1% improvement on the most challenging L-4 sequences. When the model is replaced with the even larger E-3 model, the IoU score on L-4 sequences improves by a remarkable 10.9%, clearly demonstrating that larger models are indeed better at handling difficult cases.
Figure 12: Initialize DyTrack with partially visible objects.
@@
Initializing with partially visible objects. We test DyTrack under varying degrees of target visibility in the start frame. The results are presented Fig. 12. In Figure 12 (a), the target deer exhibits relatively clear and visible features, such as its head and the front part of its body. Consequently, DyTrack can accurately and stably track the target in subsequent frames. In Figure 12 (b), we further reduce the visible area, showing only part of the deer’s head and antlers in the start frame. DyTrack demonstrates strong robustness, successfully tracking the target in subsequent frames. This indicates that DyTrack can extract effective target cues from the start frame, even when a limited portion of the target is visible. To increase the difficulty further, we tested a scenario where only a small portion of the deer’s face is visible in the initial frame, and the target exhibits camouflage effects due to its similar color and texture to the tree trunk (Figure 12 (c)). Surprisingly, DyTrack is still able to track the target in most cases, especially in the early time of the video sequence. Initializing with a partially visible object poses a challenge for trackers. In the future, DyTrack could further enhance its robustness by incorporating techniques such as multiple template and template updating.
@@ VI Conclusion
This work proposes DyTrack, a dynamic transformer tracking framework for efficient object tracking. Different with existing efficient tracking methods that primarily focus on designing lightweight modules, DyTrack works with an instance-specific reasoning fashion that learns to assign various inference routes for different frames, reducing the computations for easy scenarios. We further introduce feature recycling and target-aware self-distillation to efficiency and precision, respectively. Extensive experiments demonstrate that DyTrack achieves a superior speed-precision trade-off compared to both high-performance and high-speed methods.